OpenStudy (anonymous):

a group of five students are selected from a class of 12 students. what is the probability that jack nd jill are both in the group

OpenStudy (anonymous):

@satellite73 help?

OpenStudy (kropot72):

This problem is sampling without replacement. The hypergeometric distribution applies. Shall I take you through it?

OpenStudy (kropot72):

\[P(Jack + Jill)=\frac{\left(\begin{matrix}2 \\ 2\end{matrix}\right)\left(\begin{matrix}10 \\ 3\end{matrix}\right)}{\left(\begin{matrix}12 \\ 5\end{matrix}\right)}=\frac{\left(\begin{matrix}10 \\ 3\end{matrix}\right)}{\left(\begin{matrix}12 \\ 5\end{matrix}\right)}\]